# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

### v2

@@ 5.3.3 Algorithm Computational Complexity Analysis
During training, our algorithm largely follows the SAC framework for policy optimization. However, unlike standard SAC, our algorithm intermittently replaces the policy network’s action outputs with those generated by an external LLM based on specific conditions. As a result, the overall training complexity consists of two components: the standard SAC update steps and the LLM inference overhead incurred during action selection. Due to each SAC training step including policy network updates with complexity $O(d_{\pi})$, Q-network updates with complexity $O(d_{Q})$, and target network updates with negligible computational cost, the training computation complexity of SAC is $O(n\cdot(d_{\pi}+2d_{Q})),$ where $n$ is the batch size [52]. Assuming the LLM is queried for $p\cdot n$ samples per batch, the training computation complexity of the LLM is $O(p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}),$ where $L_{\text{LLM}}$ is the number of tokens and $d_{\text{LLM}}$ is the LLM hidden layer size [54]. Therefore, the training computation complexity of our algorithm is $O((1-p)\cdot n\cdot(d_{\pi}+2d_{Q})+p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}).$ Despite the additional overhead introduced by occasional LLM queries, the overall computation complexity remains comparable to standard SAC, and the selective use of LLM ensures that the efficiency of the training process is largely preserved, offering a favorable trade-off between performance and cost.
@@ 6.1 Experiment Setting
$N_{GE}$ [11, 15] $v_{\max}$ 10 m/s $\mathcal{N}$ $10^{-6}$
$DV_{GE_{i}}$ [1, 3] $v_{\text{limit}}$ 5 m/s PR 20 m
$TP_{GE,i}$ 0.01 W $E_{\text{total}}$ $1\times 10^{6}$ J $d_{\min},d_{\text{safe}}$ 5 m, 15 m
$N_{U^{\prime}}$ [3, 7] $E_{\text{limit}}$ $8\times 10^{5}$ J $\tau$ 3.6
@@ 6.1.1 Environment Parameters
The experimental environment is simplified to a two-dimensional space with $\mathbb{A}$ of 500 m in both length and width. $\mathbb{T}\mathbb{A}$ and $\mathbb{L}\mathbb{A}$ are square regions measuring $25\times 25$ m, located at the top-left and bottom-right corners of $\mathbb{A}$, respectively. Given the urban deployment scenario, the environment includes rectangular ${\mathbb{R}}_{NFZ}$}, ${\mathbb{R}}_{RZ}$}, and ${\mathbb{R}}_{BZ}$} with the range of [50, 200] m for both length and width. The number of such regions are randomly chosen from the range [1, 3]. The weights assigned to the components of the reward function are set to [8, 0.4, 0.4, 0.4, 0.6, 0.4, 0.6, 1, 0.01][8,\ 0.4,\ 0.4,\ 0.4,\ 0.6,\ 0.4,\ 0.6,\ 1,\ 0.01] and can be flexibly adjusted to reflect different task priorities. Other environmental parameters are detailed in Table III.
@@ 6.1.2 Algorithm Parameters
For the algorithm, hyperparameters used by the proposed algorithm and other baselines, including SAC+LLM, SAC, Proximal Policy Optimization (PPO) [55], and DDPG [56], are shown in Table IV. For the LLM used in our algorithm and other baselines, including LLM and SAC+LLM, we choose the DeepSeek-R1, which has a total of 672B parameters and 37B activated parameters [49].
@@ 6.2 Evaluation Metrics
  • Data collection rate (DCR): This metric quantifies the average ratio of successfully collected data by the DCU relative to the total target data volume in $N_{\text{total}}$ data collection tasks:
@@
  • Collision rate (CR): This metric quantifies the ratio of data collection tasks in which the DCU collides with OUs, $\mathbb{R}_{BZ}$, or flies into $\mathbb{R}_{NFZ}$:
@@
  • Successful landing rate (SLR): This metric quantifies the ratio of data collection tasks in which the DCU can return to $\mathbb{LA}$ before depleting its energy:
@@
  • Regulatory violation rate (RVR): This metric quantifies the ratio of tasks in which the DCU exceeds speed limits within $\mathbb{R}_{RZ}$:
@@ 6.3 Experiment Results
In this section, we present a comparative analysis between the proposed algorithm  [^1] The source code is available at: https://github.com/BJTU-STIC/DRL-LLM and several baseline methods in terms of the DCU flight trajectories and the evaluation metrics mentioned above. The parameter settings of these baselines are introduced in Section 6.1.2. The baselines are as follows:
  • Heuristic method: A rule-based strategy is used to determine the actions of the DCU for data acquisition and obstacle avoidance. The DCU sequentially collects data from the nearest GE until all data is acquired. If a static obstacle such as $\mathbb{R}_{NFZ}$ or $\mathbb{R}_{BZ}$ is encountered, the DCU performs detouring maneuvers. For dynamic obstacles like OUs, it either hovers in place or retreats to avoid collision until the OU moves away. To limit the noise impact when flying over $\mathbb{R}_{RZ}$, the DCU must reduce its speed below $v_{\text{limit}}$ in accordance with regulatory requirements [6]. Furthermore, if the remaining energy is only sufficient for returning to $\mathbb{LA}$, the DCU will immediately head back.
  • LLM: An LLM, DeepSeek-R1, is used in real-time to decide the actions of the DCU. We provide the LLM with the state information within a 50 m radius around the DCU, as well as the positions of the two nearest GEs. In addition, we also provide the LLM with $E_{\text{limit}}$ of the DCU.
@@
  • SAC: A pure deep reinforcement learning approach using the SAC without LLM integration.
@@
  • PPO: A representative on-policy reinforcement learning method.
@@
  • DDPG: A classic off-policy actor-critic method designed for continuous control.
@@
Figure 3: Training reward curves for the proposed algorithm and baseline algorithms over 4000 episodes, including SAC, PPO, DDPG, and SAC+LLM variants with different invocation probabilities.
(a) (a1)
(b) (a2)
(c) (a3)
(d) (a4)
(e) (b1)
(f) (b2)
(g) (b3)
(h) (b4)
(i) (c1)
(j) (c2)
(k) (c3)
(l) (c4)
(m) (d1)
(n) (d2)
(o) (d3)
(p) (d4)
(q) (e1)
(r) (e2)
(s) (e3)
(t) (e4)
Figure 4: Representative snapshots of DCU flight trajectories during data collection under different algorithms. (a1-a4): Heuristic; (b1-b4): LLM; (c1-c4): SAC+LLM (0.3); (d1-d4): SAC; (e1-e4): Ours. The trajectories illustrate obstacle avoidance, data collection behavior, and regulation compliance. PPO, DDPG, and SAC+LLM (1.0) are excluded for clarity, as their policies either fail to converge or produce unstable behaviors with limited comparative value. Obstacles (NFZs, BZs, RZs, and OUs), and GEs are highlighted in each subplot to illustrate the algorithms’ compliance and avoidance behavior.
6.3.1 Flight Trajectories about Baselines and Our Algorithm
Before presenting the flight trajectories, we first provide the training reward curves for the baselines related to DRL and the proposed algorithm. As illustrated in Fig. 3, these curves reflect the convergence behavior of each algorithm over 4000 training episodes. The proposed algorithm achieves the highest and most stable cumulative rewards, demonstrating superior policy quality and learning efficiency. Both SAC and PPO exhibit clear convergence trends, with PPO showing slightly slower but stable improvement. DDPG also improves steadily but converges to a lower reward plateau, suggesting limited effectiveness in handling the complex constraints of this environment. In contrast, the SAC+LLM variants with randomly invoked LLMs (at probabilities 0.3, 0.4, 0.5, and 1.0) exhibit significantly lower and more unstable reward trajectories. Notably, the SAC with random LLM 1.0 variant, where the LLM selects all actions, fails to converge effectively, highlighting the detrimental impact of fully replacing policy learning with LLM reasoning during training. This further supports our core design: selectively invoking the LLM in safety-critical contexts ensures that high-level reasoning complements rather than disrupts the DRL policy learning process. Overall, these results validate that our hybrid algorithm balances learning efficiency and decision safety, outperforming both standalone DRL methods and indiscriminate LLM integration strategies.
Fig. 4 illustrates segments of the DCU’s flight trajectories during data collection under both the proposed algorithm and the baselines. We highlight four typical stages: static obstacle avoidance, data collection paths, dynamic obstacle avoidance, and return to $\mathbb{LA}$. While PPO, DDPG, and SAC+LLM (0.4, 0.5, 1.0) are included in the training reward comparison (Fig. 3) to evaluate learning performance, their flight trajectories are not shown in the subsequent visualizations. This is because these methods either fail to converge to stable policies or yield suboptimal control behaviors that do not contribute additional insights beyond those already observed from SAC and SAC+LLM (0.3) variants. To maintain clarity and focus, we present representative trajectory comparisons only for the most informative and competitive baselines. Figs. 4(a1) and (b1) illustrate that both the heuristic and LLM approaches predominantly follow a “nearest GE” strategy for data collection. However, the LLM approach exhibits more efficient trajectory planning, demonstrating a tendency to select shorter and more obstacle-aware trajectories. Notably, Fig. 4(b2) indicates that the LLM does not rigidly adhere to the nearest GE principle; instead, it appears to balance the trade-off between obstacle avoidance and data acquisition efficiency. Despite this, as shown in Figs. 4(a4) and (b4), adherence to the nearest GE strategy tends to increase the overall flight distance. As a result, both the heuristic and the LLM approaches often encounter a scenario where the DCU reaches its energy limit before completing the data collection task, thereby necessitating an early return to $\mathbb{LA}$. Fig. 4(c) presents the performance of SAC+LLM, wherein the LLM is stochastically invoked during training to determine the DCU’s actions. This integration does not effectively enhance SAC’s decision-making capability and, in fact, leads to performance degradation. As observed in Figs. 4(c1) and (c4), this method results in suboptimal obstacle avoidance and frequently prevents the DCU from successfully returning to $\mathbb{LA}$, often exhausting its energy mid-flight. Figs. 4(d) and (e) depict the flight trajectories generated by the SAC algorithm and the proposed algorithm, respectively. From Figs. 4(d1) and (e1), it can be observed that SAC adopts a more conservative obstacle avoidance strategy, which inadvertently increases flight distance. Figs. 4(d2) and (e2) show that SAC may overlook distant GEs, resulting in reduced data acquisition. Furthermore, Figs. 4(d3) and (e3) indicate that SAC tends to omit data collection opportunities in the presence of obstacles, adversely affecting overall data completeness. Finally, Figs. 4(d4) and (e4) reveal that SAC may opt to traverse residential areas directly. Empirical testing suggests that SAC occasionally exceeds speed limits within these zones, thereby compromising compliance with operational regulations.
@@ 6.3.2 Metrics about Baselines and Our Algorithm
The experimental results in Figs. 5–10 comprehensively assess five critical performance metrics, DCR, CR, SLR, RVR, and ECR, under varying numbers of OUs, GEs, NFZs, BZs, and RZs. We analyzed various algorithms for the requirements: robust obstacle avoidance (R2), compliance (R3), and economy (R4). This further demonstrates how the algorithm in this paper addresses Challenges 1–3 mentioned in Section 1.
Fig. 5 presents the performance of different algorithms under increasing numbers of OUs, representing dynamic obstacles. Our algorithm consistently outperforms all baselines, maintaining the highest DCR and the lowest CR across all conditions, while also achieving the most energy-efficient trajectories (lowest ECR). Notably, while SAC performs reasonably well, it suffers from rising CR as $N_{U^{\prime}}$ increases. SAC+LLM achieves high DCR but exhibits extremely poor robustness, with CR exceeding 93%, indicating that the naive LLM integration without appropriate control undermines safety. DDPG and PPO offer moderate DCR improvements over heuristic and LLM methods, but their CRs remain higher and ECRs less efficient than ours. Notably, the standalone LLM method maintains a low CR at first, but both its DCR and ECR remain inferior to learning-based approaches, and CR increases sharply under denser conditions. These results highlight our algorithm’s superiority in dynamic, uncertain environments (R2, Challenges 1&3), and its capability to maintain economic efficiency (R4, Challenge 2) without compromising safety.
Fig. 6 evaluates the algorithms under increasing numbers of GEs, reflecting higher task complexity. Our algorithm consistently achieves the highest DCR (DCR $\approx$ 1.0 and near-perfect SLR (SLR $\approx$ 1.0), demonstrating mission reliability. While SAC also performs well, its DCR and SLR decline slightly with complexity. DDPG and PPO achieve moderate DCR and SLR across all GE levels, outperforming heuristic and LLM methods in data collection yet trailing behind SAC and our algorithm in both reliability and safety. Their CRs remain relatively higher than our algorithm and SAC, indicating less robustness when handling complex tasks. SAC+LLM exhibits high DCR at low complexity but suffers a sharp drop in SLR (to as low as 5.63%) and persistently high collision risk, indicating poor robustness in full-mission execution. The LLM alone maintains stable SLR, but its DCR degrades significantly as targets increase, limiting its effectiveness. These results highlight our algorithm’s ability to plan efficient and reliable trajectories in increasingly complex environments, addressing Challenge 2 and Challenge 3.
Figs. 7 and  8 examine algorithmic performance under increasing numbers of NFZs and BZs, both representing static obstacles. Our algorithm maintains DCR $\approx$ 1.0 and minimal CR (CR $\approx$ 0), even as obstacle density rises, clearly demonstrating superior robust static obstacle avoidance (R2). In contrast, SAC+LLM suffers from a sharp increase in CR despite relatively high DCR, indicating that the LLM component alone is insufficient to compensate for suboptimal DRL decisions in constrained spaces. The LLM alone maintains zero CR at low obstacle counts but with consistently low DCR, reflecting limited planning capability. SAC shows moderate DCR but increasing CR (up to 15.67%), suggesting vulnerability to static obstacles. DDPG and PPO demonstrate stronger performance than SAC+LLM and LLM, achieving relatively high DCR. However, both suffer from noticeably higher CR compared to SAC and our algorithm, especially as obstacle density increases, which reflects limited robustness in handling constrained static environments. Their ECRs are moderate, indicating some economic trajectory planning but still lagging behind our algorithm. Furthermore, our algorithm consistently achieves the lowest ECR across all conditions, reinforcing its ability to optimize trajectories economically. These results affirm the effectiveness of our algorithm in addressing Challenge 1 and partially Challenge 2 (maintaining the economy).
Fig. 9 explores the algorithms’ ability to balance task completion and compliance under increasing numbers of RZs, regions where entry is permissible but potentially non-compliant. Our algorithm consistently achieves the highest DCR while maintaining low CR and RVR, showcasing strong capability in resolving Challenge 3. SAC+LLM, although maintaining high DCR, incurs both high CR and RVR, suggesting its reliance on the LLM is insufficient to enforce compliance under ambiguity. SAC alone shows moderate RVR but suffers from increasing CR. DDPG and PPO also demonstrate decent task completion with relatively high DCR. However, both algorithms exhibit noticeably higher CR and RVR compared to SAC and our algorithm, indicating their limited ability to handle ambiguous regulatory boundaries. While they outperform SAC+LLM in compliance, their safety and adherence still fall short in soft-constrained environments. In contrast, the LLM and heuristic methods avoid almost all violations (0.34% and 0%), but their performance stagnates at low DCR (74.62% and 76.92%), implying over-conservatism and suboptimal data acquisition. Overall, our algorithm uniquely achieves both mission effectiveness and regulatory adherence, excelling in environments with soft constraints.
@@
In summary, the results validate that our algorithm, by selectively engaging the LLM for obstacle avoidance, empowers the DCU to make informed and adaptive decisions in complex, dynamic environments. This targeted integration enables our algorithm to uniquely balance the three core demands of the low-altitude economy: R2, R3, and R4. By meeting these requirements, our algorithm effectively addresses the corresponding Challenges 1–3. No other baseline achieves comparable comprehensive performance across these dimensions. Thus, our algorithm stands out as a practical solution for urban low-altitude data collection scenarios.
Figure 5: The DCR, CR, and ECR about baselines and our algorithm when $N_{U^{\prime}}$ increases. The proposed algorithm consistently achieves the highest DCR with almost zero CR and lowest ECR, demonstrating superior efficiency and safety.
Figure 6: The DCR, CR, and SLR about baselines and our algorithm when $N_{GE}$ increases. Our algorithm maintains a stable and high DCR and CR of almost zero, while other baselines show declining performance and increased collisions. This highlights our algorithm’s robustness in complex environments.
Figure 7: The DCR, CR, and ECR about baselines and our algorithm when $N_{NFZ}$ increases. While other algorithms suffer from reduced data collection or increased collisions, our algorithm sustains near-perfect DCR and near zero CR, with competitive energy consumption.
Figure 8: The DCR, CR, and ECR about baselines and our algorithm when $N_{BZ}$ increases. The proposed algorithm preserves a high DCR and almost avoids collisions entirely, while other baselines experience noticeable performance degradation.
Figure 9: The DCR, CR, and RVR about baselines and our algorithm when $N_{RZ}$ increases. Our algorithm maintains full compliance (RVR 0), with high data collection efficiency and no collisions, while other baselines fails to adapt to regulatory constraints.

### v3

@@ 5.3.3 Algorithm Computational Complexity Analysis
During training, our algorithm largely follows the SAC framework for policy optimization. However, unlike standard SAC, our algorithm intermittently replaces the policy network’s action outputs with those generated by an external LLM based on specific conditions. As a result, the overall training complexity consists of two components: the standard SAC update steps and the LLM inference overhead incurred during action selection. Due to each SAC training step including policy network updates with complexity $O(d_{\pi})$, Q-network updates with complexity $O(d_{Q})$, and target network updates with negligible computational cost, the training computation complexity of SAC is $O(n\cdot(d_{\pi}+2d_{Q})),$ where $n$ is the batch size [47]. Assuming the LLM is queried for $p\cdot n$ samples per batch, the training computation complexity of the LLM is $O(p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}),$ where $L_{\text{LLM}}$ is the number of tokens and $d_{\text{LLM}}$ is the LLM hidden layer size [48]. Therefore, the training computation complexity of our algorithm is $O((1-p)\cdot n\cdot(d_{\pi}+2d_{Q})+p\cdot n\cdot L_{\text{LLM}}\cdot d_{\text{LLM}}^{2}).$
It is important to clarify that the LLM is employed only during the offline training phase to assist the DRL process, rather than during real-time deployment. Specifically, the LLM provides high-level semantic guidance to the SAC agent, helping it select more informative and regulation-compliant actions during exploration, thereby accelerating policy convergence and improving training stability. Once the SAC policy network is fully trained, the LLM is completely removed from the control loop. During online execution, the UAV relies solely on the lightweight SAC policy for real-time trajectory planning and obstacle avoidance. This design ensures that the runtime control algorithm operates efficiently and achieves collision-free navigation without invoking any LLM inference, thereby fully guaranteeing real-time performance and computational feasibility.
@@ 6.1 Experiment Setting
$N_{\mathrm{GE}}$ [11, 15] $v_{\max}$ 10 m/s $\mathcal{N}$ $10^{-6}$
$DV_{\mathrm{GE}_{i}}$ [1, 3] $v_{\text{limit}}$ 5 m/s PR 20 m
$TP_{{\mathrm{GE}}_{i}}$ 0.01 W $E_{\text{total}}$ $1\times 10^{6}$ J $d_{\min},d_{\text{safe}}$ 5 m, 15 m
$N_{\mathrm{OU}}$ [3, 7] $E_{\text{limit}}$ $8\times 10^{5}$ J $\tau$ 3.6
@@ 6.1.1 Environment Parameters
The experimental environment is simplified to a two-dimensional space with $\mathbb{A}$ of 500 m in both length and width. $\mathbb{T}\mathbb{A}$ and $\mathbb{L}\mathbb{A}$ are square regions measuring $25\times 25$ m, located at the top-left and bottom-right corners of $\mathbb{A}$, respectively. Given the urban deployment scenario, the environment includes rectangular ${\mathbb{R}}_{\mathrm{NFZ}}$}, ${\mathbb{R}}_{\mathrm{RZ}}$}, and ${\mathbb{R}}_{\mathrm{BZ}}$} with the range of [50, 200] m for both length and width. The number of such regions are randomly chosen from the range [1, 3]. The weights assigned to the components of the reward function are set to [8, 0.4, 0.4, 0.4, 0.6, 0.4, 0.6, 1, 0.01][8,\ 0.4,\ 0.4,\ 0.4,\ 0.6,\ 0.4,\ 0.6,\ 1,\ 0.01] and can be flexibly adjusted to reflect different task priorities. Other environmental parameters are detailed in Table III.
@@ 6.1.2 Algorithm Parameters
For the algorithm, hyperparameters used by the proposed algorithm and other baselines, including SAC+LLM, SAC, Proximal Policy Optimization (PPO) [49], DDPG [50], and Constrained Policy Optimization (CPO) [53], are shown in Table IV. For the LLM used in our algorithm and other baselines, including LLM and SAC+LLM, we choose the DeepSeek-R1, which has a total of 672B parameters and 37B activated parameters [44].
@@ 6.2 Evaluation Metrics
  • Data collection rate (DCR): This metric quantifies the average ratio of successfully collected data by the DCU relative to the total target data volume in $N_{\text{total}}$ data collection:
@@
  • Collision rate (CR): This metric quantifies the ratio of trajectory planning tasks in which the DCU collides with OUs, $\mathbb{R}_{\mathrm{BZ}}$, or flies into $\mathbb{R}_{\mathrm{NFZ}}$:
@@
  • Successful landing rate (SLR): This metric quantifies the ratio of trajectory planning tasks in which the DCU can return to $\mathbb{LA}$ before depleting its energy:
@@
  • Regulatory violation rate (RVR): This metric quantifies the ratio of tasks in which the DCU exceeds speed limits within $\mathbb{R}_{\mathrm{RZ}}$:
@@ 6.3 Experiment Results
In this section, we present a comparative analysis between the proposed algorithm and several baseline methods in terms of the DCU flight trajectories and the evaluation metrics mentioned above. The parameter settings of these baselines are introduced in Section 6.1.2. The baselines are as follows:
  • Heuristic method: A greed-based policy [^1] We referenced the code from: https://github.com/Rosenkrands/sar-path-planning is used to determine the actions of the DCU for data acquisition and obstacle avoidance. If a static obstacle such as $\mathbb{R}_{\mathrm{NFZ}}$ or $\mathbb{R}_{\mathrm{BZ}}$ is encountered, the DCU performs detouring maneuvers. For dynamic obstacles like OUs, it either hovers in place or retreats to avoid collision until the OU moves away. To limit the noise impact when flying over $\mathbb{R}_{\mathrm{RZ}}$, the DCU must reduce its speed below $v_{\text{limit}}$ in accordance with regulatory requirements [7]. Furthermore, if the remaining energy is only sufficient for returning to $\mathbb{LA}$, the DCU will immediately head back.
  • LLM: An LLM, DeepSeek-R1, is used in real-time to decide the actions of the DCU. We provide the LLM with the state information within a 20 m radius around the DCU, as well as the positions and remaining data of GEs. In addition, we also provide the LLM with $E_{\text{limit}}$ of the DCU.
  • SAC+Heuristic: Similar to SAC+LLM, this variant employs the heuristic rule-based policy for obstacle avoidance instead of using the LLM. Specifically, when the DCU encounters potential collisions with OUs or static obstacles such as $\mathbb{R}_{\mathrm{NFZ}}$ or $\mathbb{R}_{\mathrm{BZ}}$, the heuristic controller overrides the SAC policy to execute avoidance or speed-reduction maneuvers according to predefined rules. This baseline serves to evaluate the effectiveness of integrating domain-specific expert knowledge into the SAC framework.
@@
  • SAC: A pure deep reinforcement learning approach using the SAC without LLM integration [^2] We referenced the code from: https://github.com/DLR-RM/stable-baselines3, which also implements PPO and DDPG, all available within the same repository.
@@
  • PPO: A representative on-policy reinforcement learning method, implemented within the same codebase as SAC, following the same architecture and parameterization.
@@
  • DDPG: A classic off-policy actor-critic method designed for continuous control, also implemented within the SAC codebase, ensuring consistency in experimental setup.
  • CPO: A constrained reinforcement learning baseline based on safety‑starter‑agents from OpenAI [^3] We referenced the code from: https://github.com/openai/safety-starter-agents. It optimizes the expected cumulative reward while enforcing a safety constraint on the discounted cumulative cost. In our implementation, the instantaneous cost $c_{t}$ is defined as a weighted combination of safety violations, including boundary collisions, OUs collisions, tall buildings collisions, and entries into NFZs. The threshold $d$ is set to 0.05, corresponding to an acceptable violation rate of 5%. The policy and value networks follow the same architecture as SAC, and the trust-region constraint on policy update is enforced with a maximum KL divergence of 0.01.
  • Existing methods: We also include two recent RL approaches [51, 52] for comparison to benchmark against state-of-the-art path planning and cooperative UAV control methods [^4] We referenced the code for Theile et al. [51] from: https://github.com/theilem/uavSim [^5] We referenced the code for Wang et al. [52] from: https://github.com/BIT-MCS/DRL-UCS-AoI-Threshold.
@@
Figure 4: Training reward curves of the proposed algorithm and all baseline methods over 4000 episodes, including SAC, PPO, DDPG, CPO, Theile et al. [51], Wang et al. [52], SAC+Heuristic, and SAC+LLM variants with different invocation probabilities.
6.3.1 Training Reward Curves about DRL Baselines and Our Algorithm
We show the training reward curves of all DRL baselines and the proposed algorithm, as illustrated in Fig. 4. Over 4000 training episodes, the proposed method achieves the highest and most stable cumulative rewards, indicating superior learning efficiency and policy quality. SAC and PPO converge well, with PPO showing slightly slower improvement, while DDPG reaches a lower reward plateau under complex constraints. The CPO algorithm exhibits stable but lower convergence due to its conservative optimization. Theile et al. [51] achieves faster and more stable convergence through equivariant ensemble regularization, whereas Wang et al. [52] converges slowly with lower final rewards. The SAC+Heuristic method shows early stability but limited long-term performance. Overall, the proposed hybrid LLM-augmented DRL algorithm achieves the best trade-off among efficiency, stability, and safety, outperforming both traditional and enhanced baselines.
Figure 5: The DCR, CR, and ECR about baselines and our algorithm when $N_{\mathrm{OU}}$ increases. The proposed algorithm consistently achieves the highest DCR with almost zero CR and lowest ECR, demonstrating superior efficiency and safety.
Figure 6: The DCR, CR, and SLR about baselines and our algorithm when $N_{\mathrm{GE}}$ increases. Our algorithm maintains a stable and high DCR and CR of almost zero, while other baselines show declining performance and increased collisions. This highlights our algorithm’s robustness in complex environments.
Figure 7: The DCR, CR, and ECR about baselines and our algorithm when $N_{\mathrm{NFZ}}$ increases. While other algorithms suffer from reduced data collection or increased collisions, our algorithm sustains near-perfect DCR and near zero CR, with competitive energy consumption.
Figure 8: The DCR, CR, and ECR about baselines and our algorithm when $N_{\mathrm{BZ}}$ increases. The proposed algorithm preserves a high DCR and almost avoids collisions entirely, while other baselines experience noticeable performance degradation.
Figure 9: The DCR, CR, and RVR about baselines and our algorithm when $N_{\mathrm{RZ}}$ increases. Our algorithm maintains full compliance (RVR 0), with high data collection efficiency and no collisions, while other baselines fails to adapt to regulatory constraints.
Figure 11: Variations of DCR, CR, ECR, SLR, and RVR under different hyperparameter settings of (a) actor LR, (b) critic LR, and (c) batch size.
Figure 12: Variations of DCR, CR, ECR, SLR, and RVR under different reward weight settings for $r_{1}$, $r_{2}$, $r_{5}$, and $r_{8}$.
Figure 13: Sensitivity analysis of the LLM invocation threshold the threshold. The figure shows the impact of varying $d_{\text{th}}\in\{5,10,15,20,25\}$ m on five key performance metrics (DCR, CR, ECR, SLR, and RVR). Excessively small thresholds (e.g., 5 m) cause delayed LLM activation, leading to higher collision rates and lower DCR, while overly large thresholds (e.g., 25 m) result in frequent LLM invocations and increased energy consumption. The threshold of 15 m achieves the best trade-off between safety, efficiency, and energy consumption, validating the choice used in the main experiments.
Figure 14: Ablation on reasoning objectives embedded in the LLM prompt. Bars compare the Complete Prompt with four variants that remove one objective at a time (without Safety, without Compliance, without Data efficiency, without Energy efficiency) across five metrics (DCR, CR, ECR, SLR, RVR).
@@ 6.3.2 Metrics about Baselines and Our Algorithm
The experimental results in Figs. 5–10 evaluate five core metrics (DCR, CR, SLR, RVR, and ECR) under varying numbers of OUs, GEs, NFZs, BZs, and RZs, verifying how our algorithm satisfies the requirements of robust obstacle avoidance (R2), compliance (R3), and economy (R4) while addressing Challenges 1–3.
Fig. 5 shows the performance under increasing OUs. Our method consistently achieves the highest DCR and the lowest CR and ECR, confirming its robustness and energy efficiency in dynamic environments. SAC and PPO converge well but suffer from rising CR, while SAC+LLM becomes unstable with CR exceeding 90%. DDPG performs moderately, and the standalone LLM and heuristic methods remain conservative with limited DCR. Among new baselines, CPO achieves slightly lower DCR but lower CR than PPO, Theile et al. [51] and Wang et al. [52] perform between SAC and PPO, and SAC+Heuristic outperforms the heuristic baseline with better safety and efficiency.
Fig. 6 presents the results under increasing GEs. Our algorithm maintains DCR $\approx$ 1.0 and SLR $\approx$ 1.0, showing excellent reliability. SAC, PPO, and DDPG exhibit moderate declines in DCR and SLR as complexity rises. SAC+LLM collapses under high complexity, while CPO reduces CR at the cost of minor DCR loss. Theile et al. [51] and Wang et al. [52] remain close to PPO, and SAC+Heuristic balances performance and safety better than the rule-based baseline. These results demonstrate the superior adaptability of our method to complex mission scenarios.
Figs. 7 and 8 examine static obstacle cases (NFZs and BZs). Our approach maintains DCR $\approx$ 1.0 and CR $\approx$ 0, achieving the best obstacle avoidance and lowest ECR. SAC and PPO perform well but show rising CR with denser obstacles, while CPO achieves slightly lower DCR yet improved safety. Theile et al. [51] and Wang et al. [52] remain moderate, and SAC+Heuristic improves both DCR and CR over the heuristic baseline. These confirm the algorithm’s effectiveness in static, constrained environments (Challenges 1–2).
Fig. 9 evaluates compliance under increasing RZs. Our method sustains the highest DCR with minimal CR and RVR, indicating excellent trade-offs between efficiency and regulatory adherence. SAC+LLM achieves high DCR but violates more frequently, while SAC, PPO, and DDPG show moderate compliance. CPO further lowers violations, Theile et al. [51] and Wang et al. [52] stay between SAC and PPO, and SAC+Heuristic enhances both safety and DCR compared to heuristic control. Overall, our algorithm uniquely balances safety, compliance, and efficiency under soft constraints (Challenge 3).
@@
In summary, the results validate that our algorithm, by selectively engaging the LLM for obstacle avoidance, empowers the DCU to make informed and adaptive decisions in complex, dynamic environments. This targeted integration enables our algorithm to uniquely balance the three core demands of the low-altitude economy: R2, R3, and R4. By meeting these requirements, our algorithm effectively addresses the corresponding Challenges 1–3. No other baseline achieves comparable comprehensive performance across these dimensions. Thus, our algorithm stands out as a practical solution for urban low-altitude data collection trajectory planning scenarios.
6.3.3 Parameters Sensitive Analysis
To comprehensively evaluate the robustness and generalization capability of the proposed algorithm, we conduct a series of sensitivity analyses covering four aspects: (1) hyperparameters of the training process, (2) weight coefficients of the reward function, (3) the invocation threshold distance for triggering the LLM-based decision process, and (4) the reasoning objectives embedded in the LLM prompt. The detailed analyses and corresponding results are presented as follows.
As shown in Fig. 11, variations in the actor and critic learning rates and batch size cause only minor fluctuations in performance. DCR and SLR stay close to 1.0, while CR and RVR remain nearly zero, indicating stable learning and strong safety. ECR varies slightly within [0.76, 0.78], showing consistent energy efficiency. Overall, the proposed model exhibits high robustness to hyperparameter changes.
To evaluate the impact of reward composition, we vary the weights of $r_{1}$, $r_{2}$, $r_{5}$, and $r_{8}$, which correspond to task efficiency, obstacle avoidance, compliance, and energy return, respectively. As shown in Fig. 12, increasing $r_{1}$ enhances DCR and SLR but slightly raises ECR and CR, while higher $r_{2}$ effectively lowers CR at a small cost of DCR. Increasing $r_{5}$ improves compliance (lower RVR) but slightly reduces efficiency, and larger $r_{8}$ values yield better energy use and landing stability with a minor DCR drop. Overall, the proposed algorithm remains stable and well-balanced under moderate reward weight variations.
As shown in Fig. 13, the variation of the LLM invocation threshold exhibits a clear trade-off among safety, efficiency, and energy consumption. When the threshold is too small (e.g., 5 m), the LLM is triggered too late, leading to occasional collisions and reduced data collection efficiency. Conversely, excessively large thresholds (e.g., 25 m) cause premature or frequent invocations, increasing energy consumption and slightly reducing stability. Moderate thresholds such as 10 m and 20 m achieve similar performance to 15 m, but the latter offers the best balance, maintaining high DCR and SLR with low CR and ECR. Therefore, $15$ m is adopted as the default setting in all subsequent experiments.
We ablate the four reasoning objectives in the LLM prompt by removing one component at a time and compare against the complete prompt shown in Fig. 3. The result is shown in Fig. 14. Removing Safety increases CR and slightly reduces DCR, confirming its primary role in risk control. Removing Compliance markedly degrades RVR despite similar CR, indicating the necessity of explicit compliance guidance. Removing Data efficiency lowers DCR and slightly changes ECR, showing that prioritizing informative targets contributes to sustained task completion. Removing Energy efficiency harms energy behavior and landing stability, demonstrating that explicit energy-aware reasoning helps maintain safe mission termination. Overall, the complete prompt provides the best balance across safety, compliance, and efficiency.
In summary, all sensitivity analyses consistently demonstrate the robustness and adaptability of the proposed algorithm under diverse parameter and configuration settings. Neither hyperparameter tuning nor moderate reward reweighting leads to noticeable performance degradation, and both the LLM invocation threshold and reasoning objectives exhibit stable behavior with interpretable effects. These results collectively validate that the proposed hybrid SAC–LLM algorithm maintains a well-balanced trade-off among safety, compliance, efficiency, and energy performance, even when the underlying parameters or decision conditions vary across different environments.